OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4

OpenStudy (anonymous):



OpenStudy (anonymous):

Opening to the left

OpenStudy (anonymous):

I dont know the rest

OpenStudy (anonymous):

a is 1/4p where the focus is (h, k + p)

OpenStudy (anonymous):

So 1/4p where focus is 0 , 0 + 0 ?

OpenStudy (anonymous):

(1/(4*0))(x-0)^2 + 0 ?

OpenStudy (michele_laino):

Please try to apply the definition of a parabola, namely:  so we can write this: \[\sqrt {{{\left( {{x_0} - 4} \right)}^2} + {{\left( {{y_0} - \left( { - 4} \right)} \right)}^2}} = \left| {{x_0} - 4} \right|\]